Grade 11General Physics

A machine gun fires 360 bullets per minute and each bullet travels with a velocity of 600m/s.If the mass of each bullet is 5 grams,find the power of the machine gun

Seema
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Number of bullets per second = 360/60 = 6
velocity = 600 m/s
mass = 5g = 0.005 kg

energy required for firing 1 bullet = 1/2 × 0.005 × 600^2 = 900J

For 6 bullets per sec, power = 900×6 J/s = 5400 W